# Poisson approximation of prime divisors of shifted primes

### v1

@@ 4. The proof of Theorem 1
By taking the implied constant in the conclusion of Theorem 1 sufficiently large, we may assume without loss of generality that $y$ satisfies (3.2). We also assume that $x$ is sufficiently large in terms of $a$. Fix $\theta$ with $\alpha<\theta<\gamma$, $\varepsilon>0$ satisfying $\theta-\varepsilon>\alpha$, and $\delta>0$ sufficiently small as a function on $\theta$. Recall that
@@
where $\pi^{*}(x)=\pi(x)-\pi(|a|)$. As noted earlier, when $g_{y}(m)=0$, we have $\Phi_{m}(x,y)=0$ and thus such $m$ do not contribute to (4.1). Now break up the set $\{m\in{\mathbb{Z}}:P^{+}(m)\leqslant y,g_{y}(m)\neq 0\}$ into five pieces (the conditions $P^{+}(m)\leqslant y$ and $g_{y}(m)\neq 0$ assumed in all):
@@
By Lemma 3.5,
@@
When $m>(x+a)/y$, $\Phi_{m}(x,y)$ is 1 if $m-a$ is prime and $m-a\leqslant x$, and zero otherwise. Start with the triangle inequality and (3.4), and use the same analysis as in the case $m\in\mathcal{M}_{3}$:
@@
$\ll\frac{\log x}{x}\#\{|a|<p\leqslant x:P^{+}(p+a)\leqslant y\}+e^{-(1-\varepsilon)u\log u},$
where we applied (3.2) in the last step, assuming that $\delta$ is sufficiently small. By Theorem 1 of [37] and standard bounds for the Dickman function $\rho(u)$ (see, e.g., [40, Ch. III.5]),
#$\#\{p\leqslant x:P^{+}(p+a)\leqslant y\}\ll\pi(x)u\rho(u)\ll\pi(x)e^{-(1-\varepsilon)u\log u}.$
@@ 5. Poisson approximation: the proof of Theorem 2
6. The transferrence principle: proof of Theorem 4
Assume Hypothesis $Z(\gamma)$ with $\gamma>\frac{1}{2}$ and that $\frac{1}{2}<\alpha<\gamma$. Define $F(n)=(f_{1}(n),\ldots,f_{m}(n))$. Let

### v3

@@ 4. The proof of Theorem 1
By taking the implied constant in the conclusion of Theorem 1 sufficiently large, we may assume without loss of generality that $y$ satisfies (3.2). Furthermore, if $y\leqslant x^{1/\log_{2}x}$, that is, $u\geqslant\log_{2}x$, then $e^{-\alpha u\log u}\ll_{a,A}(\log x)^{-A}$. Hence, the conclusion in Theorem 1 for $y<x^{1/\log_{2}x}$ follows from the conclusion for $y=x^{1/\log_{2}x}$. We therefore may assume that
(4.1) $x^{1/\log_{2}x}\leqslant y\leqslant x^{\delta}.$
We also assume that $x$ is sufficiently large in terms of $a$. Fix $\theta$ with $\alpha<\theta<\gamma$, $\varepsilon>0$ satisfying $\theta-\varepsilon>\alpha$, and $\delta>0$ sufficiently small as a function on $\theta$. Recall that
@@
where $\pi^{*}(x)=\pi(x)-\pi(|a|+1)$. As noted earlier, when $g_{y}(m)=0$, we have $\Phi_{m}(x,y)=0$ and thus such $m$ do not contribute to (4.2). Now break up the set $\{m\in{\mathbb{Z}}:P^{+}(m)\leqslant y,g_{y}(m)\neq 0\}$ into five pieces (the conditions $P^{+}(m)\leqslant y$ and $g_{y}(m)\neq 0$ assumed in all):
@@
Using Lemma 3.5 when $\ell\geqslant 1$ (and recalling (4.1)), and Mertens’ theorem when $\ell=0$, we obtain
@@
When $m>(x+a)/y$, $\Phi_{m}(x,y)$ is 1 if $m-a$ is prime larger than $|a|+1$ and $m-a\leqslant x$, and zero otherwise. Start with the triangle inequality and (3.4), and use the same analysis as in the case $m\in\mathcal{M}_{3}$:
@@
$\ll\frac{\log x}{x}\#\{|a|+1<p\leqslant x:P^{+}(p+a)\leqslant y\}+e^{-(1-\varepsilon)u\log u},$
where we applied (3.2) in the last step, assuming that $\delta$ is sufficiently small. By Theorem 1 of [37] [^1] This is stated only for $a=-1$ but the proof works for any $a$ with no essential changes. and standard bounds for the Dickman function $\rho(u)$ (see, e.g., [40, Ch. III.5]),
#$\#\{|a|+1<p\leqslant x:P^{+}(p+a)\leqslant y\}\ll\pi(x)u\rho(u)\ll\pi(x)e^{-(1-\varepsilon)u\log u}.$
@@ 5. Poisson approximation: the proof of Theorem 2
6. The transference principle: proof of Theorem 4
Assume Hypothesis $Z(\gamma)$ with $\gamma>0$ and that $0<\alpha<\gamma$. Define $F(n)=(f_{1}(n),\ldots,f_{m}(n))$. Let